OpenStudy (moonlitfate):

Solve the logarithmic equation algebraically. Approximate the result to three decimal places. (If there is no solution, enter NO SOLUTION.) log2 x + log2(x − 3) = 2

OpenStudy (anonymous):

did you attempt

OpenStudy (moonlitfate):

Yes, but it only confused me. I'm not entirely sure how to start solving the equation.

OpenStudy (moonlitfate):

I tried taking the log2 of both sides to try and get rid of the log2 on the left side of the equation...

OpenStudy (anonymous):

use LogAB=LOG A+LOG B

OpenStudy (anonymous):

Where A=2X AND B= 2(X-3)

OpenStudy (anonymous):

try it out, and let me know how it goes before I give you the solution

OpenStudy (moonlitfate):

2x^2 - 6 = 2

OpenStudy (moonlitfate):

2x^2 = 8 x^2 = 4 x = 2 or x= -2

OpenStudy (moonlitfate):

I tried 2, though and it isn't the right answer. Am I even somewhat close?

OpenStudy (anonymous):

nope, it should be 2x*(2x-6)

OpenStudy (anonymous):

assuming it is log to base 2 and NOT LOG to base 10 we have 4x^2-12x=2^2 4x^2-12x-4=0

OpenStudy (anonymous):

if it is log to base 10 WHICH I THINK IS MORE APPROPRIATE, it should be 4x^2-12x=10^2 4x^2-12x-100=0 which can be reduced to 3x^2-3x-25=0

OpenStudy (moonlitfate):

You have to raise 2 to the second power on the right side?

OpenStudy (anonymous):

yes ONLY IF IT IS LOG TO BASE 2, AND IF IT IS LOG TO BASE 10 we have 10^2

OpenStudy (moonlitfate):

Still doesn't make too much sense. :/

OpenStudy (anonymous):

well you need to solve the equation , i leave that to you, try to factorize, or use quadratic formula if applicable